# Magic Mirror on the Wall, How to Benchmark Quantum Error Correction Codes, Overall ?

### 20 Feb 2024 v2

@@ 2.2.8 Heavy Hexagon Code
On this lattice, data qubits and ancillas are strategically distributed along the hexagonal vertices and edges. A designated subset of these ancilla qubits serves as flag qubits, crucial for the detection of high-weight errors that may result from a limited number of fault events. The stabilizers tasked with the detection of X-type errors are constituted by the product of weight-two Z-type gauge operators, which collectively form the surface code stabilizers. Meanwhile, the X-type stabilizers, reminiscent of columnar operators in the Bacon-Shor code, are deduced by aggregating products of weight-four and weight-two X-type gauge operators. Fig. 8 represents the layout and stabilizer circuits of a heavy-hexagon code.
3 Benchmarking Methodology
This section outlines the benchmarking parameters, explains the rationale behind the selection of codes & parameters, and discusses the methodology used for comparative analysis.
3.1 Explanation of Benchmarking Parameters
3.1.1 Qubit Overhead
Qubit Overhead is a fundamental parameter in assessing the efficiency of a QECC. It refers to the number of physical qubits needed to encode a logical qubit; lower overhead is preferable as it implies fewer resources for error correction. However, this must be balanced against the code’s ability to correct errors, which often requires more qubits for increased distance. Therefore, analyzing the trade-off between qubit overhead and error correction capabilities is crucial in evaluating the practicality of a QECC.
3.1.2 Error Threshold
The Error Threshold is a critical benchmark for the robustness of a quantum error correction code. It represents the maximum physical error rate under which a QECC can effectively correct errors. A higher threshold indicates a more fault-tolerant code, making it more desirable for practical quantum computing where error rates are non-negligible.
3.1.3 Error Protection
Error Protection capability is the measure of a QECC’s ability to identify and correct a range of quantum errors, including bit flips, phase flips, and their combinations (Pauli errors). The ultimate goal is to design QECCs capable of detecting and correcting all possible quantum errors. A QECC’s versatility in handling diverse error types is a significant indicator of its effectiveness and practical utility in quantum computing environments.
3.1.4 Decoding
The availability and efficiency of Decoding Algorithms are essential for the practical implementation of QECCs. Decoding is the process of interpreting syndrome measurements to identify and correct errors. The more decoding algorithms a QECC is compatible with, the more flexible and adaptable it is. The performance of a QECC is heavily reliant on the efficiency of these algorithms, which should be fast and accurate to maintain the integrity of quantum information.
3.1.5 Transversal Gates
Transversal Gates play a crucial role in the implementation of quantum error correction. They facilitate operations on encoded qubits without decoding them, thereby preserving the error correction capability. The complexity and type of transversal gates that a QECC supports can significantly affect its functionality and integration with quantum circuits. Advanced techniques like lattice surgery, while complex, offer enhanced capabilities in error correction and logical operations, marking an important consideration in QECC design.
3.1.6 Scalability
Scalability is an essential factor for the practical application of QECCs in large-scale quantum computing. It refers to a QECC’s ability to maintain its error correction effectiveness as the system size increases or as it operates in environments with higher levels of noise. Scalability is crucial for the advancement of quantum computing, as it determines whether a QECC can be effectively used in more complex and realistic quantum systems.
3.1.7 Realization
The Realization of a QECC across different qubit types is a testament to its adaptability and practical relevance. The versatility of a QECC in being implemented with various qubit technologies (such as superconducting qubits, trapped ions, etc.) is vital for its applicability in diverse quantum computing architectures. This parameter is a strong indicator of a QECC’s potential for widespread adoption in the quantum computing field.
3.1.8 Complexity
The Structural and Operational Complexity of a QECC impacts its implementability and efficiency. This encompasses the intricacy of its encoding, error detection, and correction processes. A less complex QECC might be easier to implement but may offer limited error correction capabilities. Conversely, a more complex QECC might provide robust error correction but at the cost of increased resource demands and operational challenges. Balancing complexity with effectiveness and resource efficiency is key in the design and selection of QECCs for practical quantum computing.
3.2 Criteria for Selection of Codes and Parameters
@@
The eight chosen quantum error correction codes illustrate the historical and conceptual evolution of QECCs. Starting from basic concepts like the quantum repetition code to more complex structures like the heavy-hexagon code, they form a kind of ‘family tree’ that shows the development and diversification of QECCs over time. They encompass a wide array of error correction strategies, from simple error correction (quantum repetition code) to sophisticated techniques like those used in the surface and color codes. This diversity ensures a comprehensive understanding of different error correction methodologies. They vary significantly in terms of complexity and scalability. For example, the Shor code and Steane code represent more traditional approaches, whereas the toric and surface codes are geared toward large-scale quantum computing. This range allows for the assessment of QECCs across different scalability requirements. Different codes are optimized for different error models. Including a variety of codes ensures that your benchmarking study addresses a wide range of error scenarios. Some codes, like the surface code, are designed with practical implementation in mind, considering current quantum computing limitations. This inclusion allows for an assessment of QECCs not just in theory but also in their practical realization. The selection spans codes with varying qubit overheads. For instance, the quantum repetition code is simple but not efficient, whereas codes like the toric code aim to balance error correction efficacy with resource efficiency. These eight codes collectively cover the spectrum of QECCs, from basic to advanced, from theoretical to practical. This range is crucial for a comprehensive benchmarking study. While there are other QECCs, including more might lead to redundancy without significantly adding to the understanding of QECC performance. The chosen codes are sufficiently representative of the broader field. This selection allows for a comparative analysis that can highlight the strengths and weaknesses of different approaches to quantum error correction. It enables a holistic view of the current state of QECC technology.
@@
The eight selected parameters encompass all critical dimensions of QECC functionality - from basic resource requirements (like qubit overhead) to advanced operational aspects (like error protection and decoding). This wide-ranging coverage ensures no key aspect of QECC performance is overlooked. They strike a balance between the efficiency of a QECC (in terms of resource utilization and complexity) and its robustness (error threshold and protection capabilities). This balance is crucial for practical quantum computing where resources are limited and error rates are non-negligible. Parameters such as scalability and realization reflect the real-world challenges of implementing QECCs in various quantum computing environments. They ensure that the assessment of QECCs goes beyond theoretical effectiveness to include practical viability. Each parameter addresses a distinct aspect of QECC performance. There is minimal overlap between them, ensuring that each provides unique and valuable insight into the QECC’s capabilities. While these parameters are comprehensive based on the current understanding and needs of quantum computing, they also allow for flexibility. As quantum technology evolves, these criteria can be adapted or expanded to include new developments and requirements. Having a set of widely accepted parameters aids in standardizing the assessment of QECCs. This standardization is crucial for comparative analysis and for guiding improvements in QECC designs.
3.3 Approach to Comparative Analysis
The primary objective of this comparative analysis is to establish a comprehensive benchmark for QECCs. This benchmark serves as a reference point for future developments in QECCs, guiding researchers and practitioners in evaluating new codes against established parameters. The analysis acknowledges that the ‘best’ QECC depends on specific operational contexts, such as the number of qubits available, the type of errors prevalent in the system, and the practical constraints of quantum computing environments.
@@ 3.3 Approach to Comparative Analysis
Each selected QECC will be evaluated based on predefined criteria, including qubit overhead, error threshold, error protection, decoding, transversal gates, scalability, realization, and complexity. This structured approach ensures a comprehensive and uniform assessment of each code’s capabilities and limitations.
The analysis aims to highlight that while no single QECC is universally superior, each has its strengths and weaknesses. Through this structured and comprehensive approach to comparative analysis, the study will provide valuable insights into the relative merits of different QECCs. It will aid in the development of more effective, efficient, and contextually appropriate QECCs, thereby advancing the field of quantum computing. The ultimate goal is to foster an environment where QECC selection is based on an informed, nuanced understanding of each code’s performance in relation to specific operational needs and constraints.
4 Benchmarking Analysis
This section begins by examining the performance of eight distinct QECCs against a set of eight benchmarking parameters. It then progresses to discuss the significance of benchmarking variations within existing QECCs, emphasizing the importance of such assessments.
4.1 Comparative Benchmarking of QECCs
@@
Parameters Quantum Repetition Code Shor Code Steane Code Toric Code Surface Code Bacon-Shor Code 3D Color Code Heavy-Hexagon Code Qubit Overhead Error Threshold Error Protection Only half of all bit flip errors. Two qubits, all error type Detects errors on 2 qubits, corrects 1 All Pauli errors All Pauli errors All Pauli errors All Pauli errors All Pauli errors Decoding Automation, ML CNOT/ H Gates CNOT/ H Gates MWPM, Tensor, 10+ Algo MWPM, Tensor, 10+ Algo Statistical Mapping, Check operators Restriction, MaxSAT, 3+ Algo MWPM, ML, Neural Network, 3+ Algo Transversal Gates None None Clifford Gates TPG/Lattice Surgery Lattice Surgery Teleportation Lattice Surgery Lattice Surgery Scalability No No No Yes Yes Yes Yes Yes Realization Superconducting, Trapped ion, etc Trapped ion, Optical systems Trapped ion, Rydberg atom Superconducting, Rydberg atom arrays Superconducting, Rydberg atom arrays, Ising anyons Trapped ion None Superconducting Complexity Very Low Low Low Very High High Medium Extremely High Extremely High
@@
Table 1 shows the parameter analysis for all quantum error correction codes used in this study.
@@
In terms of qubit overhead, it is expressed as a function of the distance variable, . Specifically, the Shor and Steane codes have fixed qubit overheads of and , respectively. The repetition code’s overhead increases linearly with . The overhead for the toric and surface codes grows at a rate proportional to , while for the Bacon-Shor code, it escalates in proportion to . The 3D color code and the heavy-hexagon code exhibit qubit overheads of and , respectively. Figure 9 (left) illustrates the rise in the number of qubits corresponding to increasing values of . For instance, a Bacon-Shor code at distance demands over qubits, highlighting the significance of qubit overhead as a key benchmarking metric.
@@
The threshold error rate is a critical metric for quantum error correction codes (QECCs), indicating the highest level of physical errors they can effectively correct. This makes it a key consideration in evaluating the performance of QECCs. The threshold error rates for various codes like the repetition, Shor, Steane, toric, surface, Bacon-Shor, 3D color, and heavy-hexagon codes are [16], [65], [67], [41], [16, 37], [5], [14, 21] and [8] respectively, as depicted in Fig. 9 (right). While the repetition code exhibits the highest threshold, it is important to note its limitation to only detecting bit-flip errors, rendering it impractical for broader applications. Thus, finding a balance in these characteristics is crucial for effective quantum error correction.
@@
Different QECCs have different types and ranges of errors they can correct. The repetition code detects bit flip errors on qubits but does not detect any phase-flip errors [60]. Conversely, the Shor code detects two-qubit errors or corrects an arbitrary single-qubit error [65]. The Steane code is a distance 3 code, detecting errors on 2 qubits and correcting errors on 1 qubit [67]. More complex codes like the toric, surface, 3D color, and heavy-hexagon codes can detect and correct all Pauli errors. However, their effectiveness depends on the code’s distance, which is directly related to the qubit overhead and the complexity of implementation. The toric code on a torus is a code, encoding two qubits, while the surface code is a code, encoding only one qubit [30]. The Bacon-Shor code variant, code, has a distance [6]. Unlike the surface code, the color code can suffer from unremovable hook errors due to the specifics of its syndrome extraction circuits, requiring fault-tolerant decoders to use additional flag qubits [10]. In contrast, the heavy-hexagon code is an example of the Bacon-Shor type stabilizer [15].
@@
A decoding algorithm is essential for interpreting stabilizer syndrome values to identify the type and location of errors in a quantum error correction code (QECC). These algorithms introduce additional complexity to QECCs, and the performance of these codes is highly dependent on such algorithms. The availability of multiple decoding algorithms enhances the versatility of a code. The repetition code can use classical 2D Ising model decoders [75] or machine learning-based approaches [20]. The Shor and Steane codes employ CNOT and Hadamard gates for decoding [50]. The toric and surface codes are supported by a variety of decoding algorithms, including minimum weight perfect-matching (MWPM) [28], tensor decoders [12], union-find decoders [22], brief perfect matching [38], renormalization group methods [26], Markov-chain Monte Carlo techniques [40], cellular automata [35], neural networks [76], sliding and parallel window decoders [72], generalized belief propagation [54], among others. The Bacon-Shor code uses the mapping effect of noise to statistical mechanical models for decoding [62]. Although its check operators are few-body, stabilizer weights scale with the number of qubits, and stabilizer expectation values are obtained by taking products of gauge-operator expectation values [36]. The 3D color code may utilize various decoders such as the projection decoder [45], integer-program-based decoder [68], restriction decoder [14], cellular-automation decoder [61], or the MaxSAT-based decoder [7].
@@
The repetition, Shor, and Steane codes are not scalable across multiple qubits, in contrast to the toric, surface, Bacon-Shor, 3D color, and heavy-hexagon codes, which are scalable. The implementation complexity of these codes can be gauged from their circuit or lattice structures. The repetition code has a very low complexity. The Shor and Steane codes can be categorized as having low complexity. The surface code, with its square lattice structure, is considered highly complex. Similarly, the toric code, despite its resemblance in structure, is deemed highly complex due to its encoding of two qubits instead of one. The Bacon-Shor code falls into the medium complexity category. Finally, the lattice structures of the 3D color and heavy-hexagon codes indicate their extremely high complexity in terms of implementation.
@@
4.2 Benchmarking QECC Variations

### 05 Mar 2024 v3

@@ II-B Descriptions of QECCs Under Study
Heavy Hexagon Code: Heavy hexagon codes [39] leverage a hexagonal lattice enhanced with strategically placed ’heavy’ qubits, improving error correction efficiency through increased qubit connectivity. This design boosts error tolerance, crucial for precise syndrome measurements in identifying and correcting bit-flip (X) and phase-flip (Z) errors. The mechanism ensures the integrity of quantum information by accurately pinpointing and fixing errors. This lattice stabilizer subsystem code synthesizes elements from both Bacon-Shor and surface-code frameworks, and encodes a single logical qubit into physical qubits where denotes the code distance. This lattice is characterized by its sparse connectivity, typically not exceeding three connections per qubit, making it a robust candidate for systems with fixed-frequency transmon qubits that are vulnerable to frequency collision errors. On this lattice, data qubits and ancillas are strategically distributed along the hexagonal vertices and edges. A designated subset of these ancilla qubits serves as flag qubits, crucial for the detection of high-weight errors that may result from a limited number of fault events. The stabilizers tasked with the detection of X-type errors are constituted by the product of weight-two Z-type gauge operators, which collectively form the surface code stabilizers. Meanwhile, the X-type stabilizers, reminiscent of columnar operators in the Bacon-Shor code, are deduced by aggregating products of weight-four and weight-two X-type gauge operators. Fig. 8 represents the layout and stabilizer circuits of a heavy-hexagon code.
III Benchmarking Methodology
III-A Explanation of Benchmarking Parameters
Qubit Overhead: Qubit overhead is a fundamental parameter in assessing the efficiency of a QECC. It refers to the number of physical qubits needed to encode a logical qubit; lower overhead is preferable as it implies fewer resources for error correction. However, this must be balanced against the code’s ability to correct errors, which often requires more qubits for increased distance. Therefore, analyzing the trade-off between qubit overhead and error correction capabilities is crucial in evaluating the practicality of a QECC. Error Threshold: The error threshold is a critical benchmark for the robustness of a QECC as it represents the maximum physical error rate under which a QECC can effectively correct errors. A higher threshold indicates a more fault-tolerant code, making it more desirable for practical quantum computing where error rates are non-negligible. Error Protection: Error protection capability is the measure of a QECC’s ability to identify and correct a range of quantum errors, including bit flips, phase flips, and their combinations (Pauli errors). The ultimate goal is to design QECCs capable of detecting and correcting all possible quantum errors. A QECC’s versatility in handling diverse error types is a significant indicator of its effectiveness and practical utility in quantum computing environments. Decoding: The availability and efficiency of decoding algorithms are essential for the practical implementation of QECCs. Decoding is the process of interpreting syndrome measurements to identify and correct errors. The more decoding algorithms a QECC is compatible with, the more flexible and adaptable it is. The performance of a QECC is heavily reliant on the efficiency of these algorithms, which should be fast and accurate to maintain the integrity of quantum information. Transversal Gates: Transversal gates play a crucial role in the implementation of quantum error correction. They facilitate operations on encoded qubits without decoding them, thereby preserving the error correction capability. The complexity and type of transversal gates that a QECC supports can significantly affect its functionality and integration with quantum circuits. Advanced techniques like lattice surgery, while complex, offer enhanced capabilities in error correction and logical operations, marking an important consideration in QECC design. Scalability: Scalability is an essential factor for the practical application of QECCs in large-scale quantum computing. It refers to a QECC’s ability to maintain its error correction effectiveness as the system size increases. Scalability is crucial for the advancement of quantum computing, as it determines whether a QECC can be effectively used in more complex and realistic quantum systems. Realization: The Realization of a QECC across different qubit types is a testament to its adaptability and practical relevance. The versatility of a QECC in being implemented with various qubit technologies (such as superconducting qubits, trapped-ions, etc.) is vital for its applicability in diverse quantum computing architectures. This parameter is a strong indicator of a QECC’s potential for widespread adoption in the quantum computing field. Complexity: The Structural and Operational Complexity of a QECC impacts its implementability and efficiency. This encompasses the intricacy of its encoding, error detection, and correction processes. A less complex QECC might be easier to implement but may offer limited error correction capabilities. Conversely, a more complex QECC might provide robust error correction but at the cost of increased resource demands and operational challenges. Balancing complexity with effectiveness and resource efficiency is key in the design and selection of QECCs for practical quantum computing.
III-B Criteria for Selection of Codes and Parameters
@@
The eight chosen quantum error correction codes illustrate the historical and conceptual evolution of QECCs. Starting from basic concepts like the quantum repetition code to more complex structures like the heavy-hexagon code, they form a kind of ‘family tree’ that shows the development and diversification of QECCs over time. They encompass a wide array of error correction strategies, from simple error correction (quantum repetition code) to sophisticated techniques like those used in the surface and color codes. This diversity ensures a comprehensive understanding of different error correction methodologies. They vary significantly in terms of complexity and scalability. For example, the Shor code and Steane code represent more traditional approaches, whereas the toric and surface codes are geared toward large-scale quantum computing. This range allows for the assessment of QECCs across different scalability requirements. Different codes are optimized for different error models. Including a variety of codes ensures that the benchmarking study addresses a wide range of error scenarios. Some codes, like the surface code, are designed with practical implementation in mind, considering current quantum computing limitations. This inclusion allows for an assessment of QECCs not just in theory but also in their practical realization. The selection spans codes with varying qubit overheads. For instance, the quantum repetition code is simple but not efficient, whereas codes like the toric code aim to balance error correction efficacy with resource efficiency. These eight codes collectively cover the spectrum of QECCs, from basic to advanced, from theoretical to practical. This range is crucial for a comprehensive benchmarking study. While there are other QECCs, including more might lead to redundancy without significantly adding to the understanding of QECC performance. The chosen codes are sufficiently representative of the broader field. This selection allows for a comparative analysis that can highlight the strengths and weaknesses of different approaches to quantum error correction. It enables a holistic view of the current state of QECC technology.
@@
The eight selected parameters encompass all critical dimensions of QECC functionality - from basic resource requirements (like qubit overhead) to advanced operational aspects (like error protection and decoding). This wide-ranging coverage ensures no key aspect of QECC performance is overlooked. They strike a balance between the efficiency of a QECC (in terms of resource utilization and complexity) and its robustness (error threshold and protection capabilities). This balance is crucial for practical quantum computing where resources are limited and error rates are non-negligible. Parameters such as scalability and realization reflect the real-world challenges of implementing QECCs in various quantum computing environments. They ensure that the assessment of QECCs goes beyond theoretical effectiveness to include practical viability. Each parameter addresses a distinct aspect of QECC performance. There is minimal overlap between them, ensuring that each provides unique and valuable insight into the QECC’s capabilities. While these parameters are comprehensive based on the current understanding and needs of quantum computing, they also allow for flexibility. As quantum technology evolves, these criteria can be adapted or expanded to include new developments and requirements. Having a set of widely accepted parameters aids in standardizing the assessment of QECCs.
III-C Approach to Comparative Analysis
@@
The analysis acknowledges that the ‘best’ QECC depends on specific operational contexts, such as the number of qubits available, the type of errors prevalent in the system, and the practical constraints of quantum computing environments. Each selected QECC will be evaluated based on predefined criteria, including qubit overhead, error threshold, error protection, decoding, transversal gates, scalability, realization, and complexity. This structured approach ensures a comprehensive and uniform assessment of each code’s capabilities and limitations. The analysis aims to highlight that while no single QECC is universally superior, each has its strengths and weaknesses. Through this structured and comprehensive approach to comparative analysis, the study will provide valuable insights into the relative merits of different QECCs.
IV Benchmarking Analysis
IV-A Comparative Benchmarking of QECCs
@@
Parameters Quantum Repetition Code Shor Code Steane Code Toric Code Surface Code Bacon-Shor Code 3D Color Code Heavy-Hexagon Code Qubit Overhead Error Threshold Error Protection Only half of all bit flip errors. Two qubits, all error type Detects errors on 2 qubits, corrects 1 All Pauli errors All Pauli errors All Pauli errors All Pauli errors All Pauli errors Decoding Automation, ML CNOT/ H Gates CNOT/ H Gates MWPM, Tensor, 10+ Algo MWPM, Tensor, 10+ Algo Statistical Mapping, Check operators Restriction, MaxSAT, 3+ Algo MWPM, ML, Neural Network, 1+ Algo Transversal Gates None None Clifford Gates TPG/Lattice Surgery Lattice Surgery Teleportation Lattice Surgery Lattice Surgery Scalability No No Yes Yes Yes Yes Yes Yes Realization Superconducting, Trapped ion, etc Trapped ion, Optical systems Trapped ion, Rydberg atom Superconducting, Rydberg atom arrays Superconducting, Rydberg atom arrays, Ising anyons Trapped ion None Superconducting Complexity Very Low Low Low Very High High Medium Extremely High Extremely High
@@
Table I shows the parameter analysis for all quantum error correction codes used in this study.
@@
In terms of qubit overhead, it is expressed as a function of the distance variable, . Specifically, the Shor and Steane codes have fixed qubit overheads of and , respectively. The repetition code’s overhead increases linearly with . The overhead for the toric and surface codes grows at a rate proportional to , while for the Bacon-Shor code, it escalates in proportion to . The 3D color code and the heavy-hexagon code exhibit qubit overheads of and , respectively. Figure 9 (left) illustrates the rise in the number of qubits corresponding to increasing values of . For instance, a Bacon-Shor code at a distance demands over qubits, highlighting the significance of qubit overhead as a key benchmarking metric.
@@
The threshold error rate is a critical metric for quantum error correction codes (QECCs), indicating the highest level of physical errors they can effectively correct. The threshold error rates for various codes like the repetition, Shor, Steane, toric, surface, Bacon-Shor, 3D color, and heavy-hexagon codes are [40], [24], [26], [16], [40, 41], [42], [38, 43] and [44] respectively, as depicted in Fig. 9 (right). While the repetition code exhibits the highest threshold, it is important to note its limitation to only detecting bit-flip errors, rendering it impractical for broader applications. Thus, finding a balance in these characteristics is crucial for effective quantum error correction.
@@
Different QECCs have different types and ranges of errors they can correct. The repetition code detects bit flip errors on qubits but does not detect any phase-flip errors [45]. Conversely, the Shor code detects two-qubit errors or corrects an arbitrary single-qubit error [24]. The Steane code is a distance code, detecting errors on two qubits and correcting errors on one qubit [26]. More complex codes like the toric, surface, 3D color, and heavy-hexagon codes can detect and correct all Pauli errors. However, their effectiveness depends on the code’s distance, which is directly related to the qubit overhead and the complexity of implementation. The toric code on a torus is a code, encoding two qubits, while the surface code is a code, encoding only one qubit [46]. The Bacon-Shor code variant, code, has a distance [15]. Unlike the surface code, the color code can suffer from unremovable hook errors due to the specifics of its syndrome extraction circuits, requiring fault-tolerant decoders to use additional flag qubits [47]. In contrast, the heavy-hexagon code is an example of the Bacon-Shor type stabilizer [39].
@@
A decoding algorithm is essential for interpreting stabilizer syndrome values to identify the type and location of errors in a quantum error correction code (QECC). These algorithms introduce additional complexity to QECCs, and the performance of these codes is highly dependent on such algorithms. The availability of multiple decoding algorithms enhances the versatility of a code. The repetition code can use classical 2D Ising model decoders [48] or machine learning-based approaches [49]. The Shor and Steane codes employ CNOT and Hadamard gates for decoding [50]. The toric and surface codes are supported by a variety of decoding algorithms, including minimum weight perfect-matching (MWPM) [51], tensor decoders [52], union-find decoders [53], brief perfect matching [54], renormalization group methods [55], Markov-chain Monte Carlo techniques [56], cellular automata [57], neural networks [58], sliding and parallel window decoders [59], generalized belief propagation [60], among others. The Bacon-Shor code uses the mapping effect of noise to statistical mechanical models for decoding [61]. Although its check operators are few-body, stabilizer weights scale with the number of qubits, and stabilizer expectation values are obtained by taking products of gauge-operator expectation values [62]. The 3D color code may utilize various decoders such as the projection decoder [63], integer-program-based decoder [64], restriction decoder [38], cellular-automation decoder [65], or the MaxSAT-based decoder [66]. The heavy-hexagon code has two decoders available primarily for itself, one using machine learning [44] and the other using a neural network [67], along with MWPM.
@@
The repetition and Shor codes are not scalable across multiple qubits, in contrast to the Steane, toric, surface, Bacon-Shor, 3D color, and heavy-hexagon codes, which are scalable. The implementation complexity of these codes can be gauged from their circuit or lattice structures. The repetition code has a very low complexity. The Shor and Steane codes can be categorized as having low complexity. The surface code, with its square lattice structure, is considered highly complex. Similarly, the toric code, despite its resemblance in structure, is deemed highly complex due to its encoding of two qubits instead of one. The Bacon-Shor code falls into the medium complexity category. Finally, the lattice structures of the 3D color and heavy-hexagon codes indicate their extremely high complexity in terms of implementation.
@@
IV-B Benchmarking QECC Variations